OpenStudy (anonymous):

The equilibrium system N2O4(g) 2NO2(g) was established in a 1.00-liter vessel. Upon analysis, the following information was found: [NO2] = 0.500 M; [N2O4] = 0.0250 M. What is the value of Keq? 20. 0.050 10.0 0.10

OpenStudy (anonymous):

no one?

OpenStudy (aaronq):

Do you know how to write equilibrium expressions?

OpenStudy (aaronq):

Like this: aA + bB \(\rightleftharpoons\) cC + dD \(K_{fwd}=\dfrac{[C]^c*[D]^d}{[A]^a*[B]^b}\)

OpenStudy (anonymous):

no, and i got that one wrong :(

OpenStudy (anonymous):

that looks like it makes sense though

OpenStudy (aaronq):

okay, it's pretty easy the main focus is to arrange them like this: \(K=\dfrac{[products]}{[reactants]}\) then you assign exponents according to the stoichiometric coefficients. Give it a try

OpenStudy (anonymous):

Okay so, that makes a lot more sense... so would i then just solve the equation for my question above?

OpenStudy (aaronq):

yeah, don't forget the exponent though.

OpenStudy (anonymous):

is it .10?

OpenStudy (aaronq):

hm thats not what i got. \(\dfrac{0.5^2}{0.025}=10\)

OpenStudy (anonymous):

okay, yeah. thanks man.

OpenStudy (aaronq):

no problem, dude